OpenStudy (anonymous):

Is 9a2 − 30a + 25 a perfect square trinomial? If so, what are the factors?

OpenStudy (anonymous):

yup \[a^2-2ab+b^2=(a-b)^2\] \[(3a)^2-2(3a)(5)+(5)^2=youranswer\]

OpenStudy (anonymous):

My test has four different options I'll copy and paste them